If the perimeter of a rectangle is 164 meters and the length is 2 meters less than 5 times the width, then what are the dimensions of the rectangle?

Answer by mananth(16946)

If the perimeter of a rectangle is 164 meters and the length is 2 meters less than 5 times the width, then what are the dimensions of the rectangle?
width = x
5 tims width = 5x
2 lees
5x-2
..
length = 5x-2
2*(L+W)=perimeter
..
2(x+5x-2)=164
2(6x-2)=164
12x-4=164
add 4 to both sides
12x=168
/12
x= 14 m the width.
...
length = 5x-2
5*14-2
70-2
68 m
...
CHECK
68+68+14+14=164
m.ananth@hotmail.ca